need help with part C. thanks(Financial statement analysis) The annual sales for Salco, Inc. were $4.69 million last year. The firms end-of-year balance
Answer #1
Total Asset Turnover =Sales/Total Assets
A Sales $4,690,000
B Total Assets $1,999,000
C=A/B Total Asset Turnover 2.346173087
Total Asset Turnover 2.35 times
Operating Profit margin =Operating Profit/Sales
A Operating Profit=Net Operating Income $678,000
B Sales $4,690,000
C=A/B Operating Profit Margin 0.1445629
Operating Profit Margin 14.5%
Operating Return on Assets=Net Operating Income/Total assets
A Net Operating Income $678,000
B Total assets $1,999,000
C=A/B Operating Return on Assets 0.339169585
Operating Return on Assets 33.9%
b Total Assets after new investment $3,029,000 (1999000+1030000)
Operating Profit Margin 13.40%
Sales $4,690,000
Net Operating Income=4690000*13.4%= $628,460
New Operating Return on assets 0.207481017 (628460/3029000)
New Operating Return on assets 20.75%
c New Operating Income $628,460
Interest expense=106000+48000= $154,000
Earning Before taxes $474,460
Taxes (35%) $166,061
A Net Income $308,399
Owners equity before new investment $999,500
Addition to Owners Equity=1030000/2= $515,000
B Total Equity after new investment $1,514,500
C=A/B Return on owners equity 0.203630901
Return on owners equity 20.4%